OpenStudy (anonymous):

Simplify: x^2 - 3x- 18/ x^2 - 2x - 24 divided by 1/x+4

OpenStudy (helder_edwin):

do you mean \[ \LARGE{ \frac{\frac{x^2-3x-18}{x^2-2x-24}}{\frac{1}{x+4}}} \]

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

OK do you know how to write into factors like this \[ x^2-3x-18=(x-\quad)(x+\quad) \]

OpenStudy (anonymous):

x-6 and x+3

OpenStudy (helder_edwin):

that is very good! do the same for x^2 - 2x - 24 what do you get?

OpenStudy (anonymous):

x+4 and x-6

OpenStudy (helder_edwin):

that is great. so we have this far \[ \LARGE \frac{\frac{x^2 - 3x- 18}{x^2 - 2x - 24}}{\frac{1}{x+4}}= \frac{x^2 - 3x- 18}{x^2 - 2x - 24}\cdot\frac{1}{x+4}= \frac{(x-6)(x+3)}{(x+4)(x-6)}\cdot\frac{x+4}{1} \] can you go forward?

OpenStudy (helder_edwin):

sorry in the middle it is \[ \LARGE \frac{x^2−3x−18}{x^2−2x−24}\cdot\frac{x+4}{1} \]

OpenStudy (helder_edwin):

and on the far right (it seems you might not see it) \[ \LARGE \frac{(x-6)(x+3)}{(x+4)(x-6)}\cdot\frac{x+4}{1} \]

OpenStudy (helder_edwin):

what do you get once you simplify?

OpenStudy (helder_edwin):

i'll be right back. you should be able to do it. it is easy!

OpenStudy (helder_edwin):

so what did you get?